OpenStudy (anonymous):

4) Find the four arithmetic means between -21 and -36.

OpenStudy (anonymous):

-21, ___, ___, ___, ___, -36 You have to add the common difference, d, five times to -21 to get -36. -21 + 5d = -36 5d = -15 d = -3 So the first mean is -21 - 3 = -24 The second is -24 - 3 = -27 and so on